OpenStudy (cutegirl1000):

Which numbers are solutions to the inequality x < −7, using the replacement set {−15, −10.5, −7.2, −7, 0}? Choose all answers that are correct. A. –7 B. −7.2 C. −10.5 D. −15

OpenStudy (cutegirl1000):

all except for a is what I think

OpenStudy (cutegirl1000):

so B, C, & D

OpenStudy (cutegirl1000):

@Zeronknight

OpenStudy (.mehek.):

You are correct :)

OpenStudy (cutegirl1000):

Thanks:) can you help me with three more?

OpenStudy (.mehek.):

Sure

OpenStudy (cutegirl1000):

The solution to which inequality is shown? A. y + 3 ≥ –1 B. y + 3 > –1 C. y + 3 ≤ –1 D. y + 3 < –1 https://static.k12.com/calms_media/media/1412500_1413000/1412583/1/dc6b145dc84537ee0bec246d4cdeb4bb57fd98a1/MS_PA_130528_200025.jpg

OpenStudy (cutegirl1000):

can u help? I am confused

OpenStudy (.mehek.):

There's an open circle so it's either > or <

OpenStudy (cutegirl1000):

is it <?

OpenStudy (cutegirl1000):

actually I think the answer is B

OpenStudy (.mehek.):

< is correct

OpenStudy (cutegirl1000):

so the answer is D?

OpenStudy (cutegirl1000):

Thanks... can you help me with just 2 more?

OpenStudy (.mehek.):

Yes

OpenStudy (cutegirl1000):

Solve. https://static.k12.com/bank_packages/files/media/mathml_d5a6d6d417f521e4e2a944dfcfc542a4d3ffc77b_1.gif A. t ≤ –75 B. t ≥ –75 C. t ≤ –3 D. t ≥ –3

OpenStudy (cutegirl1000):

I think it is B

OpenStudy (.mehek.):

Multiply by -5 15 * -5 = ?

OpenStudy (cutegirl1000):

-75

OpenStudy (.mehek.):

Close, when you multiply or divide by a negative, the sign flips

OpenStudy (cutegirl1000):

what do you mean?

OpenStudy (cutegirl1000):

Oh I know that

OpenStudy (cutegirl1000):

but what did I do wrong?

OpenStudy (.mehek.):

instead of ≥ , the sign flips to ≤

OpenStudy (cutegirl1000):

Oh wait.... then it should be A

OpenStudy (.mehek.):

Yes

OpenStudy (cutegirl1000):

thanks... just one more & I will also fan you

OpenStudy (.mehek.):

Ok :)

OpenStudy (cutegirl1000):

here is a screen shot of it:)

OpenStudy (.mehek.):

for #1 add -2 + 8 = ?

OpenStudy (cutegirl1000):

+6

OpenStudy (.mehek.):

and then 6 * -2 = ?

OpenStudy (cutegirl1000):

-12

OpenStudy (.mehek.):

so x ≥ -12 it can be either B or C which one do you think?

OpenStudy (cutegirl1000):

I think it is C.

OpenStudy (.mehek.):

No :/

OpenStudy (cutegirl1000):

why not? :(

OpenStudy (.mehek.):

Notice that it's greater than -12? so -11 and on is greater

OpenStudy (cutegirl1000):

oh... x is more than or = to -12 oops

OpenStudy (cutegirl1000):

i thought x was less than or = to -12

OpenStudy (cutegirl1000):

so it is B

OpenStudy (.mehek.):

so if #1 is B then #4 is C because that's one of those sign that's less/greater than or equal to

OpenStudy (cutegirl1000):

oh ok

OpenStudy (.mehek.):

now #2 -11 + 8 = ?

OpenStudy (cutegirl1000):

-3

OpenStudy (.mehek.):

4 * -3 = ?

OpenStudy (cutegirl1000):

-12

OpenStudy (.mehek.):

so x > -12

OpenStudy (cutegirl1000):

so A

OpenStudy (cutegirl1000):

which would make #3 D

OpenStudy (.mehek.):

Yep

OpenStudy (cutegirl1000):

thanks soooo much I got 100%!!!! :D

OpenStudy (.mehek.):

:D

OpenStudy (cutegirl1000):

I wrote it

OpenStudy (cutegirl1000):

now I just have to get through 6 more math tests!!!